# Open Problems, and Future Trends

### v1

@@ II-A2 Application Stage in Industrial Production
The integration of BC in industrial applications can bring forth a multitude of benefits. One such application is in the domain of supply chain and logistics, where BC can provide a solution to the lack of traceability and transparency in traditional supply chain management. By enabling the tracking of the entire life cycle of a product, from raw materials to finished goods, BC can ensure safe and convenient management of commodity flow[25].
@@
In terms of data storage, the massive amount of multimodal data generated during industrial production processes can quickly exhaust physical storage capabilities[6]. Overreliance on centralized storage systems in the Metaverse can result in a single point of failure, leading to significant economic loss. BC’s decentralized paradigm can effectively solve the problem of single point of failure while preventing data tampering through the continuous creation of new blocks.
Furthermore, BC can promote cross-enterprise production collaboration and platform interoperability by realizing cross-chain protocols and facilitating cross-chain transactions and value flow across different devices[26]. At the enterprise level, it can enable collaboration among suppliers, manufacturers, and end-users, as well as cross-chain governance among different enterprises during the production stage.
@@ II-A3 Summary
BC can provide efficient management for the Industrial Metaverse platform through mechanisms such as non-tampering, traceability, collective maintenance, and openness and transparency[27]. However, in a BC network, each participating node must verify transactions, which can lead to latency issues due to network transmission, disk writing, and other operations. Although several studies have been proposed to optimize the network topology, design new consensus algorithms, and adopt a hierarchical structure to alleviate this problem, these solutions are limited by the size of the network and the number of nodes. Therefore, balancing decentralization, security, and scalability under the Industrial Metaverse architecture is an urgent problem that needs to be addressed for sustainable enterprise development.
II-B Digital Twin
@@
II-B1 Research Advances of DT Enabling Industrial Metaverse
@@
DT enables the digitization of physical objects, while XR technology provides the medium for interaction between the two. Therefore, in addition to DT itself, a part of the works investigate the synergy between DT and XR technologies in industrial scenarios. For instance, Yang et al.[37] propose a framework for smart crane applications based on DT and XR to improve program development efficiency and Human-computer Interaction (HCI) system usability. Coupry et al.[38] investigate options for XR technologies to improve BIM and highlight the challenges of implementing BIM-based systems. Jerov et al.[39] investigate the prospects of DT and XR applications in control systems, where they develop a DT system for XR with basic interaction mechanisms and connect it to real production scenarios. While some of the above works are limited to conceptual design and specific use cases, a collaborative approach is needed to build a system that integrates DT and XR effectively.
@@
It is noteworthy to mention that, in addition to the collaborative architecture research, the real-time simulation characteristic of DT allows for monitoring the dynamic properties of physical portals for detecting various network attacks. Gupta et al.[41] suggest synchronizing the device-generated traffic through the gateway router, utilizing Support Vector Machines (SVM) for achieving regular synchronization with DT. This framework attains an accuracy rate of over 93% in detecting malicious traffic.
II-B2 Application Stage in Industrial Production
@@
Product design stage: During the product conception stage, companies utilize DT to test the production environment, constructing a virtual model to improve production by adjusting parameters in order to reduce design defect rates and production costs, while meeting preset requirements. This process entails continuous iterative optimization of both digital and physical models, interconnecting through data flow[42].
@@
Supply chain and Operations & Maintenance (O&M): DT can optimize the production process in remote equipment O&M and supply chain management. In equipment operation and maintenance, DT provides real-time data on equipment performance, including its current condition, remaining useful life, and any potential problems or failures that may occur. In the supply chain management segment, DT creates a digital copy of the supply chain network, enabling managers to monitor and optimize their operations in real-time.
Post-commercialization services: Furthermore, the Industrial Metaverse platform integrated with DT can be used to manage the traditional industrial service phase, providing DT information on the product after it has been finished and put into the market. Both manufacturers and technical service providers can access product status information on the cloud platform, which can be used to help engineers and designers improve customer experience through customized services.
II-B3 Summary
To put it succinctly, DT enables the integration of physical and digital worlds through real-time interaction and data analysis, providing a new perspective for addressing industrial challenges[43]. By combining explicit and tacit knowledge, real-time operational data can be incorporated into traditional industrial models to build DT that aid in industrial management. However, further research is needed to consider constraints within the Metaverse[44], such as security and storage, as well as to explore integration with other enabling technologies, as current efforts in this area are still in their infancy.
II-C 6G Communication Network
@@ II-B3 Summary
The Industrial Metaverse is typically data-driven and sensitive to perception latency at the physical layer. However, sensor data at the physical layer in industrial scenarios is often complex, with characteristics such as high volume, multiscale, and high noise. Additionally, the supporting technologies involved in the Industrial Metaverse often impose high demands on communication bandwidth. For example, NFTs consume significant bandwidth during digital asset validation, and the verification of digital tag sets in BC, as well as the bandwidth requirements for transactions involving different entities[45]. Furthermore, VR systems require a bit rate of 1 Gbps to provide qualified video streams visually, and real-time feedback for remote operations also necessitates high reliability and high data connection rates.
@@
5G’s URLLC are both unable to meet the aforementioned requirements. The next-generation wireless communication technology represented by 6G can provide high throughput and low-latency guarantees for the interaction between the physical and virtual worlds’ upper and lower-layer links. This is achieved through fast transaction processing to achieve efficient management. In addition to throughput, the goal of the Industrial Metaverse is to ensure high concurrency of devices in space and time, allowing all devices and users on the platform to access and leave anytime, anywhere. This increases the demand for seamless and comprehensive connectivity provided by 6G. In this regard, 6G can provide auxiliary connections through coordination with IoT devices and achieve global coverage, while its intrinsic security technology can be used to defend against unknown security threats. Fig. 6 shows the vulnerabilities, advantages, and application stage of 6G in Industrial Metaverse.
II-C1 Research Advances of 6G Enabling Industrial Metaverse
Most existing works are optimized for 6G networks and integration with the Metaverse, and metric focus varies for different application scenarios. The performance metrics in the Industrial Metaverse include throughput, latency, reliability, Age of Information, and energy efficiency[46], among which all metrics’ weights, except latency and reliability, are deeply connected and affect joint architecture design. In this regard, Cao et al.[47] discuss several metrics’ coupling relationship in-depth and argues that communication delay, reliability, and AoI are the three most crucial metrics in an IIoT environment with heterogeneous traffic. The authors collect data and control heterogeneous traffic from the short packet transmission optimization perspective. However, their specific task division is not given, and transmission jitter’s performance loss is not negligible in the actual industrial environment.
Another approach is based on overall architecture design. MEC, an essential auxiliary mechanism of the 6G network, can significantly reduce request delay due to its proximity to the source. The devices in the factory, following the edge computing hierarchy, can include sensors, IIoT devices, and servers[12]. Sensors collect industrial data, and various IIoT device states are responsible for receiving and then converge to the cloud for unified processing. In a typical scenario, some lower-level operations in industrial production can be decentralized to the lightweight mobile end, and the administrator can control the device remotely. Various 6G mobile edge technologies exist for application to the Metaverse, including wireless communication technologies (heterogeneous radio, intelligent reflective surfaces, non-orthogonal multiple access, etc.) and new communication technologies (semantic communication, holographic-type communication, tactile communication, etc.)[13].
Aslam et al.[48] present a layered Metaverse design from the perspective of layers, with different layers connected by a data exchange layer. They then demonstrate the autonomous driving remote control case and finally discuss the challenges and propose solutions for 6G and Metaverse integration, including how to deal with wireless communication security threats, submillimeter beamforming loss, and channel interference.
Since 6G mobile authorized devices in the same production environment need to share computational memory and frequencies, unreasonable resource allocation can easily lead to network congestion, which we will cover in detail in the next section.
@@ II-C1 Research Advances of 6G Enabling Industrial Metaverse
II-C2 Application Stage in Industrial Production
Computer Vision Scenarios with Multiple Accesses: With the advent of 6G-enabled industrial IoT devices, computer vision scenarios can be significantly enhanced with seamless connections and high data transfer rates, enabling better image quality in the Metaverse. Additionally, MEC resources in a virtual 6G network can leverage deep learning techniques to process heterogeneous devices in the enterprise[45].
Product Decision and Service Phase: The product decision and service phase can also benefit from 6G, as it facilitates computation-centric communication by exchanging communication resources to ensure the required quality of service for adequate computational accuracy. This, in turn, improves decision-making efficiency and enables further automation. Moreover, 6G can simplify the supply chain network, leading to agile and adaptive service provisioning for Metaverse applications.
II-C3 Summary
The network heterogeneity of each Metaverse subplatform of manufacturers presents a significant challenge for the development of a unified communication network, requiring a common network access and control protocol. Additionally, plants that deploy 6G devices must ensure the compatibility and operability of existing network-enabled devices, which must be fine-tuned to each individual subsystem. Currently, there is a lack of available 6G-enabled IIoT architectures for deployment on Industrial Metaverse platforms, as well as a lack of corresponding standards and specifications. Although relevant reference architectures have been proposed[49], the hierarchical relationships are oversimplified and require further empirical testing, and the theoretical aspects remain unexplored.
II-D Extended Reality

### v2

@@ II-A2 Application Stage in Industrial Production
The integration of BC in industrial applications can bring forth a multitude of benefits. One such application is in the domain of supply chain and logistics. Raw materials need to circulate through different stages of production to enhance their value, and this mobility is dependent on the industrial supply chain. As an enabling technology for the Industrial Metaverse, BC can be utilized to track the entire lifecycle from raw materials to finished products, addressing issues such as the lack of traceability and low transparency in traditional supply chains[25]. Operators can monitor product flows and status changes in real-time, enhancing supply chain transparency and controllability, reducing human intervention errors, and providing a secure and efficient management model for product circulation.
@@
In terms of data storage, the massive amount of multimodal data generated during industrial production processes can quickly exhaust physical storage capabilities[5]. Overreliance on centralized storage systems in the Metaverse can result in a single point of failure, leading to significant economic loss. BC’s decentralized paradigm can effectively solve the problem of single point of failure while preventing data tampering through the continuous creation of new blocks[26].
Furthermore, BC can achieve interoperability within virtual worlds through cross-chain protocols, simplifying connections using cross-chain technology. At the device level, it enables cross-chain transactions and value flows between different devices[27]. At the enterprise level, it facilitates collaboration among suppliers, manufacturers, and end-users, as well as cross-chain governance among different companies during the production stages. This reduces the complexity of system and enterprise integration, enhances operational efficiency, and allows operators to focus more on core tasks.
@@ II-A3 Summary
BC can enable efficient management of Industrial Metaverse platforms by leveraging its characteristics of immutability, traceability, collective maintenance, and transparency, thus reducing operational risks and human errors[28]. However, in a BC network, each participating node needs to validate transactions, a process that involves network transmission, disk writing, and other operations, leading to potential delays[29]. For example, during the validation process, if network latency affects the real-time synchronization of production data and task response speed, operators may face information lags, which could lead to erroneous decisions or production interruptions. While some studies have proposed optimizing network topology, designing new consensus algorithms, and adopting hierarchical structures to alleviate this issue, such solutions are still limited by network scale and the number of nodes involved. Additionally, in some industrial scenarios, trust relationships between participants have already been established or can be guaranteed through centralized authentication systems. For example, in a private industrial network led by large enterprises or multiple partners, all participants may already trust a centralized management platform or permission system.
II-B Privacy-Preserving Computing
In the context of the Industrial Metaverse, blockchain primarily relies on its immutability and consensus mechanisms to ensure the secure operation of the platform, while PPC focuses on the protection of sensitive information during cross-platform and cross-domain data sharing and computation. Technologies such as FL and DP enable secure data sharing and computation without disclosing users’ private and sensitive information, thus ensuring the security of data for both users and enterprises[30]. Fig. 5 shows the vulnerabilities, advantages, and application stage of DT in Industrial Metaverse.
II-B1 Research Advances of PPC Enabling Industrial Metaverse
FL, as a classic distributed machine learning method in the field of PPC, effectively protects data privacy by allowing data to be stored on local devices, avoiding centralized storage and data exchange. Each participant performs local computation, model training, and updates, then aggregates model parameters to the central server instead of exchanging raw data. This enables collaborative training of machine learning models without sharing sensitive data. FL in the Metaverse has been regarded as a promising method because it ensures that raw data remains local to users while also reducing the memory and computational demands on the central server by integrating FL into the Metaverse[31].
However, despite FL’s ability to aggregate data without directly uploading raw data, malicious clients could still indirectly degrade the accuracy of the global model in the Metaverse by uploading erroneous model parameters. Sun et al.[32] point out that although some research has implemented defense mechanisms to prevent server attacks during the aggregation of global models, these defenses do not mitigate the effects of high-intensity attacks. Additionally, variations in device performance lead to differences in data feature distributions and quality. When there are significant differences in participants’ data distributions, the global model has difficulty converging and suffers from poor performance. To address this, they propose a client-based defense mechanism that combines KNN with deep neural network feature representations to facilitate the rapid improvement of local model performance. Marios et al.[33] further explore personalized FL (PFL) and proposed an adversarial poisoning attack model to evaluate PFL’s susceptibility. By utilizing Bayesian optimization for layer optimization and optimizing the hidden layers of local models based on mutual information and Kullback-Leibler divergence, they aim to reduce classification uncertainty in the global model within PFL.
It is important to note that, in addition to poisoning attacks, attackers might also analyze aggregated model parameters to infer individual participants’ training data, potentially recovering sensitive user information. Therefore, HE has become an indispensable privacy-preserving technique in FL, as it allows computations to be performed on encrypted data without decrypting it.
However, FL in the Metaverse, which handles large-scale data, requires efficient and lightweight security and privacy solutions. Security and efficiency must be considered simultaneously. While HE is a reliable security mechanism, it is not suitable for large-scale industrial data scenarios due to its computational cost. FL must focus on how to protect data and from what perspective. For example, the random response method in FL is efficient, but it does not always handle the complex data generated by Metaverse devices. Some research has proposed lightweight HE methods to address this issue. For instance, Du et al.[34] introduce a lightweight HE method based on erroneous formalism learning, where the original data is encrypted as a vector and a key exchange matrix computes the final ciphertext. However, current lightweight HE methods still face efficiency challenges when processing large-scale data and lack sufficient flexibility for diverse application scenarios.
In this context, DP serves as a complementary privacy-preserving solution. It ensures that attackers cannot infer any specific individual’s private information from the model’s output by adding noise to the data during processing and analysis. Compared to HE, DP incurs lower computational overhead and is suitable for large-scale data processing and real-time systems. However, the introduction of DP in the Metaverse applications often leads to a trade-off between privacy and utility. The key research challenge in applying DP is minimizing the loss of data utility while ensuring privacy protection. Liu et al.[35] construct a basic model for DP-based Metaverse data sharing using WGAN (Wasserstein GAN) and achieved a balance between privacy and utility by incorporating appropriate regularization and Wasserstein distance. Letafati et al.[36] consider global DP in distributed Metaverse systems, drawing inspiration from the ”invisibility mode” concept, and achieved a balance between diagnostic accuracy and loss function under different privacy levels.
As the complexity of the Metaverse environment continues to grow, large language models (LLMs), as powerful natural language processing tools, are beginning to show great potential in Industrial Metaverse. Lyu et al.[37] propose a method called Differential Private Neural Representation (DPNR), which further enhances privacy protection by discarding sensitive data. However, their focus was primarily on the inference phase, neglecting privacy issues during training. To address this, Qu et al.[38] apply local DP mechanisms to fine-tune the BERT model, balancing privacy protection with model utility. Furthermore, the MA framework proposed by Abadi et al.[39] and the privacy budget algorithm by Meiser and Mohammad provide theoretical methods to calculate the privacy cost of DP, further enhancing privacy protection effectiveness.
Figure 5: The vulnerabilities, advantages, and application stage of PPC in Industrial Metaverse.
In addition to HE and DP, other privacy-preserving technologies such as Secure Multi-Party Computation (SMPC) and Zero-Knowledge Proofs (ZKP) are also continually evolving and provide alternative solutions for data privacy protection. SMPC allows multiple participants to jointly compute without revealing their private data, while ZKP enables one party to prove the truth of a statement without disclosing the underlying data. However, their practical applications in the Industrial Metaverse are still in the early stages, with relatively few research efforts.
II-B2 Application Stage in Industrial Production
In the early stages of the Industrial Metaverse, vast amounts of data are generated by various devices, sensors, and users, which often contain sensitive information. PPC technologies, such as FL, enable multiple participants to collaboratively perform computations and training tasks without disclosing their individual data, effectively protecting data privacy.
In the process of data sharing and collaboration, privacy-preserving technologies are widely applied to prevent the leakage of private information from individual users. For instance, when model parameters are shared among multiple devices or organizations, DP adds noise to ensure that attackers cannot infer specific user data from the model’s output, thereby effectively protecting privacy. Meanwhile, zero-knowledge proofs allow one party to prove the authenticity of certain data or information without disclosing the actual content, providing additional security in the process of data sharing and validation.
II-B3 Summary
Overall, with the continuous development of privacy-preserving technologies, data sharing and computation in the Industrial Metaverse will become more secure and efficient. By integrating FL, HE, DP, and other technologies, it will be possible to achieve cross-platform and cross-domain data sharing and collaboration without compromising privacy protection. Future research may focus on further optimizing these technologies, enhancing their efficiency in large-scale and real-time data processing, while maintaining effective privacy protection and reliability.
II-C Digital Twin
@@
II-C1 Research Advances of DT Enabling Industrial Metaverse
@@
DT facilitates the digitization of physical objects, while XR technologies provide the cognitive means for interaction between the digital and physical realms. As a result, beyond the scope of DT itself, several studies explore the synergy between DT and XR technologies in industrial contexts. For example, Yang et al. [49] propose a framework for smart crane applications that integrates DT and XR to enhance program development efficiency and the usability of Human-Computer Interaction (HCI) systems. Coupry et al. [50] explore the use of XR technologies to improve Building Information Modeling (BIM) and highlight the challenges associated with implementing BIM-based systems. Jerov et al. [51] examine the potential of combining DT and XR in control systems, where they develop a DT-based XR system with fundamental cognitive interaction mechanisms, linking it to real-world production scenarios. While these studies are often confined to conceptual designs or specific use cases, there is a growing need for a collaborative approach to build integrated systems that effectively combine DT and XR.
@@
II-C2 Application Stage in Industrial Production
@@
In the product design stage, enterprises initially use DT to test the production environment and build virtual models, adjusting parameters to optimize production, reduce design defect rates, and lower costs while meeting predefined requirements. This process alleviates the burden of manual trial and error, improving design accuracy and efficiency. It also involves the iterative optimization of both digital and physical models, achieving seamless connectivity through data flow[53].
@@
In the supply chain and operations & maintenance (O&M) stages, DT can optimize production processes by facilitating remote equipment maintenance and supply chain management. In equipment maintenance, DT provides real-time performance data, including current status, remaining service life, and potential issues or malfunctions, enabling operators and maintenance personnel to identify and resolve problems early, thus reducing equipment downtime.
In supply chain management, a DT of the network is created, allowing managers to monitor and optimize operations in real-time. In the post-commercialization services phase, traditional industrial services often face issues of fragmented management and lack of unified coordination, which complicates product tracking, maintenance, and support. However, an Industrial Metaverse platform integrated with DT technology can consolidate data throughout the product lifecycle, enabling manufacturers, technical service providers, and customers to access and update product status in real-time on a unified platform. This effectively addresses the issue of fragmented management and improves service efficiency and quality.
II-C3 Summary
To put it succinctly, DT can facilitate the integration of cyber-physical systems throughout the entire PLC by enabling real-time bidirectional interactions and data-driven modeling between the physical and digital worlds. This provides a new perspective for addressing challenges in industrial scenarios[54]. By combining explicit and tacit, structured and unstructured knowledge, real-time operational data can be layered over traditional industrial models to construct industrial DTs, helping operators better understand and manage manufacturing processes. However, existing research must build upon current DT technologies to address Metaverse-specific constraints[55], such as security limitations and storage restrictions. Additionally, attention should be given to deployment strategies and integration with other enabling technologies, as research in this area remains at a relatively early stage.
II-D 5G/6G Communication Network
@@ II-C3 Summary
The Industrial Metaverse is typically data-driven and sensitive to perception delays in the physical layer. However, sensor data in industrial scenarios is often highly complex, with characteristics of high volume, multi-scale, and high noise. Additionally, the supporting technologies involved in the Industrial Metaverse often have high communication bandwidth requirements, such as the substantial bandwidth consumption needed for digital asset verification in NFTs, and the bandwidth demand for verifying digital tags and transactions between different entities in blockchain. Furthermore, VR systems require a data rate of 1Gbps to provide acceptable video streaming, and real-time feedback for remote operations necessitates high reliability and high-speed data connections.
@@
Next-generation wireless communication technologies, such as 5G/6G, can provide high throughput and low latency guarantees for the upper and lower links of interactions between the physical and virtual worlds, enabling efficient management through fast transaction processing. In addition to throughput, the Industrial Metaverse aims to support high concurrency of devices within spatial and temporal bounds, ensuring that all devices and users within the platform can access and exit at any time and from any location, which increases the demand for seamless and full-coverage connectivity provided by 5G/6G. Specifically, 5G/6G technologies already support key capabilities such as URLLC, massive machine-type communication (mMTC), enhanced mobile broadband (eMBB), and network slicing, which are critical for meeting the real-time requirements of industrial services. Additionally, the new-generation wireless communication network can provide auxiliary connectivity through IoT devices for global coverage, and its inherent security technologies can also be used to defend against unknown security threats[56].
Fig. 7 shows the vulnerabilities, advantages, and application stage of 5G/6G in Industrial Metaverse.
@@
II-D1 Research Advances of 5G/6G Enabling Industrial Metaverse
Currently, the latest version of 5G technology is already capable of supporting XR applications. Devaki et al.[57] point out that fundamental XR and VR support will be enabled by 5G-Advanced (5G-A) networks; however, a complete “Metaverse” experience cannot be envisioned until 6G is deployed. Both 5G-A and 6G offer time services with many dedicated installation-based use cases, such as wireless time-sensitive networking in the Industrial IoT and critical in-vehicle networks[58]. Most existing works focus on optimizing 5G/6G networks and their integration with the Metaverse, but the emphasis on different performance metrics varies by application scenario. In the Industrial Metaverse, key performance indicators include throughput, latency, reliability, age of information (AOI), and energy efficiency[59]. While latency and reliability are deeply coupled, the weight of other metrics also influences the joint architectural design. In this regard, Cao et al.[60] discuss several metrics’ coupling relationship in-depth and argues that communication delay, reliability, and AoI are the three most crucial metrics in an IIoT environment with heterogeneous traffic. The authors collect data and control heterogeneous traffic from the short packet transmission optimization perspective. However, their specific task division is not given, and transmission jitter’s performance loss is not negligible in the actual industrial environment.
Another approach is based on the design of an integrated architecture. MEC, as a crucial auxiliary mechanism in 5G/6G networks, can significantly reduce request latency due to its proximity to the source. In a factory setting, devices can be organized into edge computing hierarchies, including sensors, IIoT devices, and servers[61]. Sensors collect industrial data, IIoT devices receive and process various states, and data is then aggregated for centralized processing in the cloud. A typical scenario involves delegating lower-tier industrial operations to lightweight mobile devices, controlled remotely by administrators. Various 5G/6G mobile edge technologies are already being applied to the Metaverse, including wireless communication technologies (e.g., heterogeneous radios, intelligent reflective surfaces, non-orthogonal multiple access) and new communication technologies (e.g., semantic communication (SemCom), holographic communication, tactile communication)[62]. Muranaka et al.[63] propose a worker navigation system using 5G and MEC that provides services without being affected by existing industrial network infrastructure. By leveraging local 5G and MEC connections, the system can effectively avoid disruptions to the existing infrastructure. Qian et al.[64] focus on optimizing user connectivity and server computing frequency allocation in an NFT-centric, BC-enabled Metaverse. By introducing the trust cost ratio as a metric, they ensured sustained user engagement and trust. Aslam et al.[65] present a layered Metaverse design from the perspective of layers, with different layers connected by a data exchange layer. They then demonstrate the autonomous driving remote control case and finally discuss the challenges and propose solutions for 5G/6G and Metaverse integration, including how to deal with wireless communication security threats, submillimeter beamforming loss, and channel interference. Since 6G mobile authorized devices in the same production environment need to share computational memory and frequencies, unreasonable resource allocation can easily lead to network congestion, which we will cover in detail in the next section.
In Table IV, we compare 5G and 6G in the context of the Metaverse across multiple dimensions. In summary, both technologies provide critical network support for the Metaverse. 5G, with its current maturity and widespread adoption, lays the foundation for the initial development of the Metaverse. Meanwhile, 6G, with its higher bandwidth, lower latency, and enhanced security and computing capabilities, is poised to become a more promising support technology in the future, delivering seamless, real-time, and highly efficient interactive experiences within the Metaverse.
TABLE IV: Comparison of 5G and 6G in Metaverse Applications
Dimension 5G 6G
Peak Data Rate Up to 20 Gb/s 100 Gb/s - 1 Tb/s
Latency 10 ms Less than 1 ms
Network Density Supports up to 1 million devices per km2 Supports up to 10-100 million devices per km2
Bandwidth Up to 1 GHz Up to 100 GHz
Time Synchronization Capabilities Advanced but limited for real-time DT applications Enhanced for precise time-sensitive applications, including XR-driven DTs
Security Features Cryptographic algorithms Quantum-safe cryptography, enhanced security protocols
Application Areas VR/AR, IoT, immersive gaming, telemedicine 3D holography, tactile internet, XR, AI-driven DTs
Computing Paradigm Primarily cloud computing Edge computing with integration of quantum computing capabilities
Energy Efficiency Significant improvement compared to 4G 10x improvement compared to 5G
Integration with IoT and IIoT Limited edge capability for large-scale IoT integration Seamless integration with IIoT through high concurrency and low latency
II-D2 Application Stage in Industrial Production
In multi-access computer vision scenarios, industrial IoT devices supported by 5G/6G can achieve seamless connectivity and communication, with dedicated frequencies and high data transfer rates that significantly enhance image quality within the Metaverse. This enables operators to access visual information more quickly and accurately. Furthermore, heterogeneous devices within enterprises can leverage multi-access edge computing resources over virtual 5G/6G networks to process data using deep learning technologies. This greatly simplifies operators’ workloads and reduces data analysis latency and complexity.
In the product decision and service phase, 5G/6G facilitates compute-centric communication by dynamically allocating communication resources to ensure the required Quality of Service (QoS) and achieve precise computation. By enhancing quality control, decision-making efficiency is improved, and further automation is enabled. Additionally, 5G/6G can reduce the complexity of supply chain networks, thus supporting agile and adaptive service delivery for Metaverse applications. Operators can respond and adjust more promptly in dynamically changing environments.
II-D3 Summary
Firstly, in terms of time synchronization, although significant advancements have been made in 5G’s time synchronization capabilities, the scope of time services in 6G still needs to be enhanced. For example, using XR to enable real-time replication of the physical world in DTs requires precise time synchronization to drive data-based decision-making for additional AI use cases, such as predictive maintenance. Accurate mirroring of the physical environment depends critically on this capability. Additionally, due to the significant network heterogeneity among various Metaverse sub-platforms of manufacturers, Metaverse communication networks represented by 5G/6G require unified network access and control protocols. Factories deploying 5G/6G devices must also ensure the operability of existing network function devices, with implementation tailored to each individual subsystem. At present, a fully deployable 5G/6G-enabled IIoT architecture for Industrial Metaverse platforms has yet to emerge, and corresponding standards and specifications remain lacking. While some recent works have proposed related reference architectures[66], they often feature overly simplified hierarchical structures and lack empirical validation, leaving theoretical aspects unexplored.
II-E Extended Reality